OpenStudy (anonymous):

an objects motion is describe by the equation d=3sin(8pi t)-2. the displacement d is measured in meter. the time t is measured in seconds. a. what is the object position (in meter)at t=0? b. what is the object maximum displacement(in meter) from its t=0 position? c. how much time(in sec) is required for one oscillation d. what is the frequency(measured in Hz) of this oscillation

OpenStudy (anonymous):

Hint a: Plug & chug. Hint b: Think about the amplitude. Hint c: The sine function has a period of \(2\pi\), so how much does \(t\) have to increase by for \(8\pi t\) to increase by \(2\pi\)? Hint d: Frequency is the inverse of period (when measured in seconds).